# Local convergence rates for Wasserstein gradient flows and McKean-Vlasov equations with multiple stationary solutions

### v2

@@ 4 Vlasov-Fokker-Planck equation
The conditions (20), (Lip), ($x-$Lip) and (U-LSI) hold for some $\lambda,L,M,\eta>0$, and (k-NL-LSI) holds on some non-empty set $\mathcal{A}\subset\mathcal{P}_{2}(\mathbb{R}^{2d})$ for some $\overline{\eta}>0$. Furthermore there exist $\rho_{*}\in\mathcal{A}\cap\mathcal{K}_{k}$ and $\delta>0$ such that $\mathcal{B}_{\mathcal{W}_{2}}(\rho_{*},\delta)\subset\mathcal{A}$ and $\mathcal{B}_{\mathcal{W}_{2}}(\rho_{*},\delta)\cap\mathcal{K}_{k}=\{\rho_{*}\}$. The function $\mathcal{L}$ is given by (65) with $a=(3+4M^{2})^{-1}$.
@@
$\max\left(\mathcal{W}_{2}^{2}(\rho_{0},\rho_{*})\,,\,\mathcal{H}(\rho_{0}|\rho_{\infty})\,,\,\mathcal{F}_{k}(\rho_{0})-\mathcal{F}_{k}(\rho_{*})\right)\\ \leqslant Ce^{-ct}\min\left(\mathcal{W}_{2}^{2}(\rho_{0},\rho_{*})\,,\,\mathcal{H}(\rho_{0}|\rho_{\infty})\,,\,\mathcal{F}_{k}(\rho_{0})-\mathcal{F}_{k}(\rho_{*})\right)\,.$ (69)
@@ 5.1 Settings and results
When the solution of the non-linear McKean-Vlasov equations converges to some $\rho_{*}$, under suitable conditions (in particular of regularity), a quantitative weak convergence at rate $1/N$ for the empirical distribution of the corresponding system of $N$ interacting particles is shown in [21, Theorem 3.1] to hold over times of order $N^{p}$ for any $p$. This requires the initial condition to start close to $\rho_{*}$ in the total variation sense. However, thanks to Proposition 6, if the initial condition is close to $\rho_{*}$ in $\mathcal{W}_{2}$, then it becomes close in the total variation sense after a time $1$.
@@ 5.2 Proofs
Similarly to Proposition 6, this follows from a Girsanov transform, see [30, Theorem 2.3] or [14, Lemma 5.3].
@@
Thanks to [24, Theorem 1], $\alpha(N)\leqslant CN^{-2/d}$ for some $C$ (but possibly for some particular $\rho_{*}$ it may vanish faster).
@@
for all $t\geqslant 1$ for some $C^{\prime}>0$. Conclusion follows by using that $t\mapsto\mathcal{F}^{N}(\rho_{t}^{N})$ is non-decreasing (distinguishing whether $t$ is above or under $\ln N/(2C)$) and applying [24, Theorem 1] to bound $\alpha(N)$. ∎
@@
The proof is the same as the proof of Proposition 21. Proposition 23 still holds in the kinetic case (with $\rho_{*}\in\mathcal{K}_{k}$; this is [14, Lemma 5.3]), and so does Lemma 24 (replacing $\mathcal{F}$ by $\mathcal{F}_{k}$, with the same proof, noticing that the kinetic energy is linear and thus satisfies (76) with $\lambda^{\prime}=0$). We use Theorem 20 instead of Theorem 8. ∎

### v4

@@ 4 Vlasov-Fokker-Planck equation
The conditions (26), (Lip), ($x-$Lip) and (U-LSI) hold for some $\lambda,L,M,\eta>0$, and (k-NL-LSI) holds on some non-empty set $\mathcal{A}\subset\mathcal{P}_{2}(\mathbb{R}^{2d})$ for some $\overline{\eta}>0$. Furthermore there exist $\rho_{*}\in\mathcal{A}\cap\mathcal{K}_{k}$ and $\delta>0$ such that $\mathcal{B}_{\mathcal{W}_{2}}(\rho_{*},\delta)\subset\mathcal{A}$ and $\mathcal{B}_{\mathcal{W}_{2}}(\rho_{*},\delta)\cap\mathcal{K}_{k}=\{\rho_{*}\}$. The function $\mathcal{L}$ is given by (71) with $a=\frac{1}{2}\left(1+2M^{2}+2\frac{L^{2}}{\sigma^{4}}\right)^{-1}$.
@@
$\max\left(\mathcal{W}_{2}^{2}(\rho_{t},\rho_{*})\,,\,\mathcal{H}(\rho_{t}|\rho_{\infty})\,,\,\mathcal{F}_{k}(\rho_{t})-\mathcal{F}_{k}(\rho_{*})\right)\\ \leqslant Ce^{-ct}\min\left(\mathcal{W}_{2}^{2}(\rho_{0},\rho_{*})\,,\,\mathcal{H}(\rho_{0}|\rho_{\infty})\,,\,\mathcal{F}_{k}(\rho_{0})-\mathcal{F}_{k}(\rho_{*})\right)\,.$ (80)
@@ 5.1 Settings and results
When the solution of the non-linear McKean-Vlasov equations converges to some $\rho_{*}$, under suitable conditions (in particular of regularity), a quantitative weak convergence at rate $1/N$ for the empirical distribution of the corresponding system of $N$ interacting particles is shown in [22, Theorem 3.1] to hold over times of order $N^{p}$ for any $p$. This requires the initial condition to start close to $\rho_{*}$ in the total variation sense. However, thanks to Proposition 6, if the initial condition is close to $\rho_{*}$ in $\mathcal{W}_{2}$, then it becomes close in the total variation sense after a time $1$.
@@ 5.2 Proofs
Similarly to Proposition 6, this follows from a Girsanov transform, see [32, Theorem 2.3] or [15, Lemma 5.3].
@@
Thanks to [26, Theorem 1], $\alpha(N)\leqslant CN^{-2/d}$ for some $C$ (but possibly for some particular $\rho_{*}$ it may vanish faster).
@@
for all $t\geqslant 1$ for some $C^{\prime}>0$. Conclusion follows by using that $t\mapsto\mathcal{F}^{N}(\rho_{t}^{N})$ is non-decreasing (distinguishing whether $t$ is above or under $\ln N/(2C)$) and applying [26, Theorem 1] to bound $\alpha(N)$. ∎
@@
The proof is the same as the proof of Proposition 21. Proposition 23 still holds in the kinetic case (with $\rho_{*}\in\mathcal{K}_{k}$; this is [15, Lemma 5.3]), and so does Lemma 24 (replacing $\mathcal{F}$ by $\mathcal{F}_{k}$, with the same proof, noticing that the kinetic energy is linear and thus satisfies (87) with $\lambda^{\prime}=0$). We use Theorem 20 instead of Theorem 8. ∎